# Directed and Undirected Vertex Connectivity Problems are Equivalent for Dense Graphs

### v1

@@ 3.2 A Reduction from Directed SSVC to Undirected SSVC
Given an algorithm ${\mathcal{A}}_{0}$ solving single-source minimum vertex cut on undirected graphs, there is an algorithm that solves the single-source minimum vertex cut problem on directed graphs, making one call to ${\mathcal{A}}_{0}$ with input of $O(n^{2})$ edges and requiring additional $O(n^{2})$ time. Moreover, the reduction holds even when we restrict both inputs to unweighted graphs.
We first describe the construction and its property.
Construction.
We are given a weighted directed graph $G=(V,E,w)$ and a source $s$. Define $A=V-N^{out}_{G}[s]$ and $\forall x\in A,w^{\prime}(x)=w(x)$, and define $A^{\prime}=\{u^{\prime}\mid u\in A\}$ as a copy of $A$ with the same node weights, i.e., $\forall x^{\prime}\in A^{\prime},w^{\prime}(x^{\prime})=w(x)$.
We define the skeleton $H=(V^{\prime},E_{H},w^{\prime})$ where
$V^{\prime}$ $=V\cup A^{\prime}$
$E_{H}$ $=E_{1}\cup E_{2}\cup E_{3}\mbox{ where }$
$E_{1}$ $=\{(s,v)\mid v\in A^{\prime}\cup N^{out}_{G}(s)\},\mbox{ connecting }s\mbox{ to all vertices but }A,$
$E_{2}$ $=\{(u,u^{\prime})\mid u\in A\},\mbox{a perfect matching between }A\mbox{ and }A^{\prime}$
$E_{3}$ $=\{(u,v)\mid u,v\in A\},\mbox{a clique inside }A.$
Note that the skeleton depends on the fixed set $N^{out}_{G}(s)$. To complete the construction, we add the remaining edges from $G$ to $H$.
We define a weighted undirected graph $G^{\prime}=(V^{\prime},E^{\prime}):=H\cup E_{4}\cup E_{5}$ where
$E_{4}=$ $\{(u,v^{\prime})\mid u,v\in A,(v,u)\in E\},$
$E_{5}=$ $\{(u,v)\mid u\in A,v\in N^{out}_{G}(s),(v,u)\in E\}.$
In $E_{4}$, for every edge $(v,u)\in E$ where $u,v\in A$, we lift its tail to $A^{\prime}$, i.e., we add $(u,v^{\prime})$ to $E_{4}$. In $E_{5}$, for every edge $(v,u)\in E$ going out of $N_{G}^{out}(s)$, we add $(u,v)$ to $E_{5}$. We define $F_{4}$ and $F_{5}$ as the set of edges in $G$ that are mapped to $E_{4}$ and $E_{5}$, respectively. This completes the construction, see Figure 1 for illustration.
Figure 1: (Left) The skeleton graph $H$. (Right) The complete construction. The set $R\subseteq A$ is denoted as a red box.
Before we describe the reduction, let us observe useful properties of $G^{\prime}$. The role of $E_{1}$ is to ensure the following proposition.
Proposition 3.4.
For every vertex cut $(L^{\prime},S^{\prime},R^{\prime})$ in $G^{\prime}$ if $s\in L^{\prime}$ then $R^{\prime}\subseteq A$
@@
Suppose $R^{\prime}$ contains some node $x\not\in A$. Thus, $s$ does not have an edge to $x$ by the definition of a vertex cut. However, by $E_{1}$, $s$ has an edge to all but $A$. In particular, $s$ has an edge to $x$, a contradiction.∎
The crux of the construction is to ensure the following property.
@@
For all $\emptyset\neq T\subseteq A,$ $w(N_{G^{\prime}}(T))=w(N_{G}^{in}(T))+w(A).$
@@
The set $N_{G}^{in}(T)$ is equal to the following set. First, we start from $N_{G^{\prime}}(T)$ in $G^{\prime}$. Then, we drop $E_{2}\cup E_{3}$ from $G^{\prime}$ (removing all neighbors of $T$ from the skeleton $H$). The remaining neighbors of $T$ use $E_{4}\cup E_{5}$. In this case, we modify those edges in $E_{4}\cup E_{5}$ by undoing the lifting in $E_{4}$ and orienting the edges in $E_{5}$ appropriately, so that they correspond to the edges in $F_{4}\cup F_{5}$ in $G$. After this process, the total reduction in the weight of $N_{G^{\prime}}(T)$ is $w(N_{H}(T))=w(A)$. Hence, $w(N_{G}^{in}(T))=w(N_{G^{\prime}}(T))-w(N_{H}(T))$, which is equivalent to $w(N_{G^{\prime}}(T))=w(N_{G}^{in}(T))+w(N_{H}(T))=w(N_{G}^{in}(T))+w(A).$
This allows us to relate the $s$-vertex mincuts between $G$ and $G^{\prime}$.
Lemma 3.6.
$\kappa_{G^{\prime}}(s)=\kappa_{G}(s)+w(A)$.
@@
We break the proof into two directions.
  • $\kappa_{G^{\prime}}(s)\geq\kappa_{G}(s)+w(A)$. Let $(L^{\prime},S^{\prime},R^{\prime})$ be an $s$-vertex mincut in $G^{\prime}$. Since $s\in L^{\prime}$, Proposition 3.4 implies that $R^{\prime}\subseteq A$. Therefore,
    $\kappa_{G^{\prime}}(s)=w(N_{G^{\prime}}(R^{\prime}))=w(N_{G}^{in}(R^{\prime}))+w(A)\geq\kappa_{G}(s)+w(A)$
    The second equality follows from Lemma 3.5.
  • $\kappa_{G^{\prime}}(s)\leq\kappa_{G}(s)+w(A)$. Let $(L,S,R)$ be an $s$-vertex mincut in $G$. By design, $R\subseteq A$. Therefore,
    $\kappa_{G}(s)+w(A)=w(N_{G}^{in}(R))+w(A)=w(N_{G^{\prime}}(R))\geq\kappa_{G^{\prime}}(s)$
    The second equality follows from Lemma 3.5. ∎
We are ready to prove Lemma 3.3.
Proof of Lemma 3.3.
Given $G$ and $s$, we construct the aforementioned graph $G^{\prime}$. We call ${\mathcal{A}}_{0}$ on $G^{\prime}$ with source $s$. Suppose the output is $(L^{\prime},S^{\prime},R^{\prime})$ with $s\in L^{\prime}$. We return the vertex cut $N^{in}_{G}(R^{\prime})$. Note that $N^{in}_{G}(R^{\prime})$ is well-defined because $R^{\prime}\subseteq A$ by Proposition 3.4.
Correctness.
We prove that this is a vertex mincut in $G$ separating $s$ from $R$. Let $(L,S,R)$ be a vertex cut in $G$ where $R=R^{\prime},S=N^{in}_{G}(R^{\prime}),L=V(G)-(S\cup R)$. We first prove that $s\in L$. Indeed, by design, there is no edge from $s$ to $A$ in $G$. Since $R\subseteq A$, there is no edge from $s$ to $R$. It remains to prove $w(S)=\kappa_{G}(s).$ By Lemmas 3.5 and 3.6, $w(S)=w(S^{\prime})-w(A)=(\kappa_{G}(s)+w(A))-w(A)=\kappa_{G}(s).$
Time Complexity.
The graph $G^{\prime}$ can be constructed in $O(n^{2})$ time and the size of the graph is $O(n^{2})$. When $G$ is unweighted, $G^{\prime}$ is also unweighted. ∎
3.3 A Reduction from Undirected SSVC to Undirected VC
Lemma 3.7.
If there is an algorithm $A_{0}$ solving minimum vertex cut on undirected graphs, then there is an algorithm solving single-source minimum vertex cut on undirected graphs making one call to ${\mathcal{A}}_{0}$ with input $O(m)$ edges and additional $O(m)$ time. Moreover, the reduction holds even when we restrict both inputs to unweighted graphs.
@@ 3.3 A Reduction from Undirected SSVC to Undirected VC
Given a single-source minimum vertex cut instance $G,s$, suppose $G$ is a weighted graph, we replace $s$ by a node with weight $W=\sum_{v\in V}w(v)$, then run ${\mathcal{A}}_{0}$ on $G$. The correctness and complexity are obvious, since $S$ cannot be in included in a minimum vertex cut.
If $G$ is an unweighted graph, we replace $s$ by a clique of size $\delta+1$ (the minimum degree of $G$) denoted by $C$, and connect all neighbors of $s$ to all vertices in the clique to get $G^{\prime}$. Let $(L,S,R)$ be a minimum vertex cut of $G^{\prime}$, firstly, $S$ cannot contain all vertices in $C$, as otherwise $S$ is larger than $\delta$. WLOG let $L$ be the vertex set that contain vertices in $C$. Then we have $C\cup N_{G}(s)\subseteq L\cup S$ according to the definition of $G^{\prime}$. Thus, we can remove all vertices in $C\cap S$ from $S$ to get $S^{\prime}$, and $S^{\prime}$ remains a vertex cut not containing $s$. In other words, if $S$ is minimum, then $S$ cannot contain any vertex in $C$. So calling ${\mathcal{A}}_{0}$ on $G^{\prime}$ suffices to find the single-source minimum vertex cut. The size of $G^{\prime}$ is $O(m)$ since there are at most $O(\delta\cdot n)$ newly added edges. ∎

### v2

@@ 3.2 Global minimum vertex cut
$\kappa_{G^{\prime}}=\kappa_{G}+w(V)$.
@@
Let $(L^{\prime},S^{\prime},R^{\prime})$ be a minimum vertex cut in $G^{\prime}$. By Proposition 3.3, we may assume without loss of generality that $L^{\prime}\subseteq V_{\text{out}}$ and $R^{\prime}\subseteq V_{\text{in}}$. Put $L=\{v\in V:v_{\text{out}}\in L^{\prime}\}$ and $R=\{v\in V:v_{\text{in}}\in R^{\prime}\}$; then $L$ and $R$ are nonempty and disjoint, and there is no edge $(u,v)\in E$ such that $u\in L$ and $v\in R$. It follows that $w^{\prime}(N_{G^{\prime}}(R^{\prime}))=w(N_{G}^{\text{in}}(R))+w(V)$ by Lemma 3.4. Since $(L^{\prime},S^{\prime},R^{\prime})$ is a minimum vertex cut and all vertex weights are positive, we have $S^{\prime}=N_{G^{\prime}}(R^{\prime})$, so the weight of the cut is $w^{\prime}(S^{\prime})=w(N_{G}^{\text{in}}(R))+w(V)$. Therefore, $(L\cup X,N_{G}^{\text{in}}(R),R)$, where $X=V\setminus(L\cup R\cup N_{G}^{\text{in}}(R))$, is a directed vertex cut in $G$ with weight $w(N_{G}^{\text{in}}(R))$. This concludes the first direction $\kappa_{G^{\prime}}\geq\kappa_{G}+w(V)$.
To prove the second direction, let $(L,S,R)$ be a minimum directed vertex cut in $G$. Let $L_{\text{out}}=\{v_{\text{out}}:v\in L\}$, $R_{\text{in}}=\{v_{\text{in}}:v\in R\}$. As $(L,S,R)$ is a directed vertex cut in $G$, the sets $L$ and $R$ are nonempty and disjoint, and there are no edges $(u,v)\in E$ such that $u\in L$ and $v\in R$. It follows from the definition of $G^{\prime}$ that there are no edges between $L_{\text{out}}$ and $R_{\text{in}}$, and the two sets are nonempty and disjoint. Let $X=V_{\text{out}}\setminus(L_{\text{out}}\cup N_{G^{\prime}}(R_{\text{in}}))$ and observe that there are no edges between $X$ and $R_{\text{in}}$. Then $(L_{\text{out}}\cup X,N_{G^{\prime}}(R_{\text{in}}),R_{\text{in}})$ is a vertex cut in $G^{\prime}$. It has weight $w(N_{G^{\prime}}(R_{\text{in}}))=w(N_{G}^{\text{in}}(R))+w(V)$ by Lemma 3.4, which gives us an upper bound on the minimum vertex cut size $\kappa_{G^{\prime}}$ in $G^{\prime}$. By the minimality of $(L,S,R)$, we have $S=N_{G}^{\text{in}}(R)$, so $\kappa_{G}=w(N_{G}^{\text{in}}(R))$ and hence $\kappa_{G^{\prime}}\leq\kappa_{G}+w(V)$. ∎
3.3 All-pairs minimum vertex cut
For the all-pairs case, we proceed analogously to the global case. We use the following notation. For distinct $s,t\in V$, let $\kappa_{G}(s,t)$ denote the weight of the minimum weighted directed $s$-$t$-cut in $G$. For distinct $s^{\prime},t^{\prime}\in V^{\prime}$, let $\kappa_{G^{\prime}}(s^{\prime},t^{\prime})$ denote the weight of the minimum weighted undirected $s^{\prime}$-$t^{\prime}$-cut in $G^{\prime}$. Note that these values are only defined if $(s,t)\notin E$ and $\{s^{\prime},t^{\prime}\}\notin E^{\prime}$, respectively; otherwise we can treat them as $\infty$. Then $\kappa_{G}(s,t)$ is embedded in $G^{\prime}$ as follows.
@@ 3.3 All-pairs minimum vertex cut
Let $s,t\in V$ such that $s\neq t$ and $(s,t)\notin E$. Then $\kappa_{G}(s,t)=\kappa_{G^{\prime}}(s_{\text{out}},t_{\text{in}})-w(V)$.
@@
Given that $s\neq t$ and there is no edge $(s,t)\in E$, there is also no edge $\{s_{\text{out}},t_{\text{in}}\}\in E^{\prime}$, hence there exists as minimum vertex cut in $G^{\prime}$ separating $s_{\text{out}}$ from $t_{\text{in}}$.
Let $(L^{\prime},S^{\prime},R^{\prime})$ be a minimum vertex cut in $G^{\prime}$ separating $s_{\text{out}}$ from $t_{\text{in}}$. Following Proposition 3.3, we may assume without loss of generality that $s_{\text{out}}\in L^{\prime}\subseteq V_{\text{out}}$ and $t_{\text{in}}\in R^{\prime}\subseteq V_{\text{in}}$. Put $L=\{v\in V:v_{\text{out}}\in L^{\prime}\}$ and $R=\{v\in V:v_{\text{in}}\in R^{\prime}\}$; then $L$ and $R$ are disjoint, $s\in L$, $t\in R$, and there is no edge $(u,v)\in E$ such that $u\in L$ and $v\in R$. It follows that $w^{\prime}(N_{G^{\prime}}(R^{\prime}))=w(N_{G}^{\text{in}}(R))+w(V)$ by Lemma 3.4. By the minimality of $(L^{\prime},S^{\prime},R^{\prime})$ and as all vertex weights are positive, we have $S^{\prime}=N_{G^{\prime}}(R^{\prime})$, so the weight of the cut is $w^{\prime}(S^{\prime})=w(N_{G}^{\text{in}}(R))+w(V)$. Therefore, $(L\cup X,N_{G}^{\text{in}}(R),R)$, where $X=V\setminus(L\cup R\cup N_{G}^{\text{in}}(R))$, is a directed vertex cut separating $s$ and $t$ in $G$ with weight $w(N_{G}^{\text{in}}(R))$. This concludes the first direction $\kappa_{G^{\prime}}(s_{\text{out}},t_{\text{in}})\geq\kappa_{G}(s,t)+w(V)$
To prove the second direction, let $(L,S,R)$ be a minimum directed vertex cut separating $s\in L$ from $t\in R$ in $G$. Let $L_{\text{out}}=\{v_{\text{out}}:v\in L\}$, $R_{\text{in}}=\{v_{\text{in}}:v\in R\}$. As $(L,S,R)$ is a directed vertex cut in $G$, the sets $L$ and $R$ are nonempty and disjoint, and there are no edges $(u,v)\in E$ such that $u\in L$ and $v\in R$. It follows from the definition of $G^{\prime}$ that $s_{\text{out}}\in L_{\text{out}}$, $t_{\text{in}}\in R_{\text{in}}$, there are no edges between $L_{\text{out}}$ and $R_{\text{in}}$, and the two sets are disjoint. Let $X=V_{\text{out}}\setminus(L_{\text{out}}\cup N_{G^{\prime}}(R_{\text{in}}))$ and observe that there are no edges between $X$ and $R_{\text{in}}$ in $G^{\prime}$. Then $(L_{\text{out}}\cup X,N_{G^{\prime}}(R_{\text{in}}),R_{\text{in}})$ is a vertex cut in $G^{\prime}$ separating $s_{\text{out}}\in L_{\text{out}}$ from $t_{\text{in}}\in R_{\text{in}}$. It has weight $w(N_{G^{\prime}}(R_{\text{in}}))=w(N_{G}^{\text{in}}(R))+w(V)$ by Lemma 3.4, which gives us an upper bound on $\kappa_{G^{\prime}}(s_{\text{out}},t_{\text{in}})$ in $G^{\prime}$. By the minimality of $(L,S,R)$, we have $S=N_{G}^{\text{in}}(R)$, so $\kappa_{G}(s,t)=w(N_{G}^{\text{in}}(R))$ and hence $\kappa_{G^{\prime}}(s_{\text{out}},t_{\text{in}})\leq\kappa_{G}(s,t)+w(V)$. ∎
3.4 Single-source/sink minimum vertex cut
Recall that, in addition to the input graph $G=(V,E)$, we are given a source vertex $s$ or a sink vertex $t$. We denote the single-source and single-sink connectivity by $\kappa_{G}(s,\ast)$ and $\kappa_{G}(\ast,t)$, respectively. As single-source and single-sink cuts in the undirected case are equivalent, we simply write $\kappa_{G^{\prime}}(v)$ for $v\in V^{\prime}$.
Corollary 3.7.
$\kappa_{G}(s,\ast)=\kappa_{G^{\prime}}(s_{\text{out}})-w(V)$ and $\kappa_{G}(\ast,t)=\kappa_{G^{\prime}}(t_{\text{in}})-w(V)$.
@@ 3.4 Single-source/sink minimum vertex cut
Using the fact that $V_{\text{out}}$ and $V_{\text{in}}$ are cliques and Lemma 3.6, we have
$\kappa_{G^{\prime}}(s_{\text{out}})$ $=\min_{v\in V}\kappa_{G^{\prime}}(s_{\text{out}},v_{\text{in}})=\min_{v\in V}\kappa_{G}(s,v)+w(V)=\kappa_{G}(s,\ast)+w(V)\ \text{ and }$
$\kappa_{G^{\prime}}(t_{\text{in}})$ $=\min_{v\in V}\kappa_{G^{\prime}}(v_{\text{out}},t_{\text{in}})=\min_{v\in V}\kappa_{G}(v,t)+w(V)=\kappa_{G}(\ast,t)+w(V).\qed$
3.5 Steiner minimum vertex cut
Recall that, in addition to the input graph $G=(V,E)$, we are given a terminal set $T$. Let $T^{\prime}=T_{\text{out}}\cup T_{\text{in}}$. We show that the undirected Steiner minimum vertex cut solution in $G^{\prime}$ with terminal set $T^{\prime}$ corresponds to a directed optimal solution.
Corollary 3.8.
$\min_{s,t\in T}\kappa_{G}(s,t)=\min_{s^{\prime},t^{\prime}\in T}\kappa_{G^{\prime}}(s^{\prime},t^{\prime})-w(V)$.
@@ 3.5 Steiner minimum vertex cut
Let $s^{\prime},t^{\prime}\in T^{\prime}$ be arbitrary. If $s^{\prime},t^{\prime}\in V_{\text{in}}$ or $s^{\prime},t^{\prime}\in V_{\text{out}}$, there is a clique edge between $s^{\prime}$ and $t^{\prime}$, so $\kappa_{G^{\prime}}(s^{\prime},t^{\prime})=\infty$. Therefore, we may without loss of generality (as $G^{\prime}$ is undirected) assume that $s^{\prime}\in V_{\text{out}}\cap T^{\prime}=T_{\text{out}}$ and $t^{\prime}\in V_{\text{in}}\cap T^{\prime}=T_{\text{in}}$. Then $s^{\prime}=s_{\text{out}}$ and $t^{\prime}=t_{\text{in}}$ for some $s,t\in T$. We may assume $s\neq t$ and $(s,t)\notin E$, otherwise we again have $\kappa_{G}(s,t)=\kappa_{G^{\prime}}(s^{\prime},t^{\prime})=\infty$. Then by Lemma 3.6, we have $\kappa_{G}(s,t)=\kappa_{G^{\prime}}(s^{\prime},t^{\prime})-w(V)$, and the claim follows. ∎
3.6 Combining Together
We are ready to prove Theorem 3.1.
Proof of Theorem 3.1.
By construction, $G^{\prime}$ has $2n$ vertices and $2\binom{n}{2}+n+m\leq O(n^{2})$ edges and the terminal set $T^{\prime}$ for the Steiner variant has size $2|T|$. By Lemmas 3.5, 3.6 and Corollaries 3.7, 3.8, the solutions in $G^{\prime}$ for global, all-pairs, single-source/sink, and Steiner variants of vertex cuts correspond to directed solutions in $G$. Moreover, given an undirected vertex cut $(L_{\text{out}},S^{\prime},R_{\text{in}})$ in $G^{\prime}$, we can efficiently extract a directed vertex cut $(L,S,R)$ in $G$, where $S=V\setminus(L\cup R)$ and $w(S)=w(S^{\prime})-w(V)$, ∎
4 Concluding Remarks
One can notice that our reduction from directed to undirected graph has a loss in boosting the number of edges to $O(n^{2})$, which is not a good reduction when the graph is moderately sparse. It is a great open problem to prove that directed vertex connectivity can be solved in almost linear time for any sparsity, not necessarily done by a black-box reduction.
Acknowledgment
This research was partially supported by the ETH Zürich Foundation, Dr. Max Rössler, and the Walter Haefner Foundation.